# spacing=nonfrench

### 20 Mar 2026 v1

@@ 5.1 Software System Verification via ITP
Formal methods have successfully verified critical software systems including operating system kernels, hypervisors, file systems and compilers. Interactive theorem proving (ITP) provides strong formal assurance by allowing engineers to encode specifications and build proofs in ITP systems such as Rocq or Isabelle/HOL. Notable examples include CertiKOS (gu2016certikos, ), which employs a compositional refinement framework to verify a layered OS kernel; CompCert leroy2016compcert , an optimizing C compiler whose translation passes are all verified to preserve semantics; FSCQ (fscq, ), a crash-safe file system proven correct under all failure scenarios using Crash Hoare Logic; seKVM hypervisor (sekvm, ), a reduced KVM core with enforced confidentiality and integrity of guest memory; seL4 klein2014comprehensive , an OS microkernel which remains the most comprehensive and influential milestone for formal system verification. These efforts demonstrate that ITP enables high-assurance proofs for critical systems, albeit at a significant engineering cost.
@@ 5.2 Automated Theorem Proving
Automated theorem proving (ATP) has long served as a benchmark for machine reasoning. Classical ATP systems rely on symbolic methods, such as resolution-based first-order provers (e.g., Vampire (riazanov2002design, ), E (schulz2019faster, )) and SMT solvers (e.g., Z3 (de2008z3, ), CVC5 (barbosa2022cvc5, )) with built-in theories. These solvers are also integrated into interactive theorem provers via hammers (e.g., Sledgehammer (bohme2010sledgehammer, ) in Isabelle), enabling automated discharge of proof goals (czajka2018hammer, ). Related to ATP, several works (e.g., Ironclad Apps (hawblitzel2014ironclad, ) and Verve (yang2011safe, )) leverage verification-aware languages (e.g., Dafny (leino2010dafny, ) and Boogie (barnett2005boogie, )) to prove memory and control-flow safety of system components, including kernels and device drivers.
@@
In recent years, machine learning has been increasingly applied to ATP. Early efforts focused on premise selection and tactic prediction using classical learning methods (e.g., Holophrasm (whalen2016holophrasm, ), TacticToe (gauthier2021tactictoe, ), Proverbot9001 (sanchez2020proverbot, ), Diva (first2022diversity, )), showing improved success rates in Metamath, HOL4, and Coq. More recently, large language models (LLMs) have been utilized to generate complete proofs or sketches (jiang2022draft, ; cao2025reviving, ), achieving possible superhuman results on mathematical problems (trinh2024solving, ). Prior attempts to extend these capabilities to software verification include Baldur (first2023baldur, ), PALM (lu2024proof, ), and Guided Proof Search (prasad2024guided, ). However, they remain limited to textbook-scale problems, which are insufficient to address challenges from real-world systems-level verification projects such as seL4.
@@
Complementary to whole-proof generation, the proof-step prediction paradigm harnesses LLMs to propose the next tactic or lemma, guided by proof-state feedback. Systems such as GPT-f (polu2020generative, ), LISA (jiang2021lisa, ), HTPS (lample2022hypertree, ) and LeanDojo (yang2023leandojo, ) demonstrate the potential of this class of methods across Metamath, Isabelle and Lean. However, this line of work primarily targets singleton mathematical problems and therefore does not scale to system-level software verification. Our method fills this gap by integrating a fine-tuned LLM for proof search with symbolic techniques, specifically, proof-step revision and proof-state filtering, making large-scale verification both effective and efficient.
@@
To further enhance reliability, hybrid approaches are proposed to incorporate symbolic provers into LLM pipelines. A series of works, e.g., Thor (jiang2022thor, ), ProofAug (liu2025proofaug, ) and Strat2Rocq (fang2025proof, ), exemplify this by coupling LLM-generated reasoning with ATPs or hammers. Yet, ATP-based verification fundamentally trades expressiveness for automation, limiting scalability in system-level contexts. In contrast, our method integrates LLMs with an ITP workflow, invoking ATPs only in the final stage to discharge the remaining proof obligations.
@@ 6 Limitations and Future Work
Rigid rule-driven step revision. The proof step revision modules heavily rely on heuristic tactic-premise recombination and lightweight premise retrieval. These mechanisms, though helpful, are imperfect and may introduce misleading candidates. Future improvements may include learning the relevance of premises directly from the proof state or integrating stronger semantic filters.
@@ 7 Conclusion
In this paper, we have presented a neuro-symbolic proof-generation framework that automates proof search for large-scale software verification projects in interactive theorem provers. Instead of attempting whole-proof synthesis, our approach performs a best-first tree search over fine-grained proof states, repeatedly querying a lightweight, fine-tuned LLM for the next proof step. This paradigm not only substantially improves data efficiency, but also enables a seamless integration of symbolic reasoning tools with LLM-based step generation and scoring, forming an effective and efficient pipeline for proof-step construction as well as proof-state filtering and ranking. The experiments on the seL4 verification corpus show our approach significantly outperform prior baselines. Notably, our method succeeds on a considerably higher proportion of multi-step, nontrivial theorems, addressing a long-standing limitation of LLM-based provers. Additional experiments on AFP projects and translated program-verification benchmarks further showcase strong cross-project generalizability.
Our work provides compelling evidence that proof-step–centric neuro-symbolic search can serve as a practical path toward scalable and automated software verification at the systems level, e.g. operating systems kernels, network architecture, and databases. Although there is still a long way to go, we are making promising progress towards developing AI-enabled, rigorous engineering methods for real-world computer systems.
Availability
The tool and all experimental data are available at the following private link, [^4] https://figshare.com/s/da4d1d995a9aa64eeadf?file=60276236 and a prebuilt Docker image containing the seL4 verification environment and all required dependencies will also be released.

### 07 May 2026 v2

@@ 6.1 Software System Verification via ITP
Formal methods have successfully verified critical software systems including operating system kernels, hypervisors, file systems and compilers. Interactive theorem proving (ITP) provides strong formal assurance by allowing engineers to encode specifications and build proofs in ITP systems such as Rocq or Isabelle/HOL. Notable examples include CertiKOS (gu2016certikos), which employs a compositional refinement framework to verify a layered OS kernel; CompCert (leroy2016compcert), an optimizing C compiler whose translation passes are all verified to preserve semantics; FSCQ (fscq), a crash-safe file system proven correct under all failure scenarios using Crash Hoare Logic; seKVM hypervisor (sekvm), a reduced KVM core with enforced confidentiality and integrity of guest memory; seL4 (klein2014comprehensive), an OS microkernel with machine-checked proofs of functional correctness and security properties. These efforts demonstrate that ITP enables high-assurance proofs for critical systems, albeit at a significant engineering cost.
@@ 6.2 Automated Theorem Proving
Automated theorem proving (ATP) has long served as a benchmark for machine reasoning. Classical ATP systems rely on symbolic methods, such as resolution-based first-order provers (e.g., Vampire (riazanov2002design), E (schulz2019faster)) and SMT solvers (e.g., Z3 (de2008z3), CVC5 (barbosa2022cvc5)) with built-in theories. These solvers are also integrated into interactive theorem provers via hammers (e.g., Sledgehammer (bohme2010sledgehammer) in Isabelle), enabling automated discharge of proof goals (czajka2018hammer). Related to ATP, several works (e.g., Ironclad Apps (hawblitzel2014ironclad) and Verve (yang2011safe)) leverage verification-aware languages (e.g., Dafny (leino2010dafny) and Boogie (barnett2005boogie)) to prove memory and control-flow safety of system components, including kernels and device drivers.
@@
In recent years, machine learning has been increasingly applied to ATP (li2024survey). Early efforts focused on premise selection and tactic prediction using classical learning methods (e.g., Holophrasm (whalen2016holophrasm), TacticToe (gauthier2021tactictoe), Proverbot9001 (sanchez2020proverbot), Diva (first2022diversity)), showing improved success rates in Metamath, HOL4, and Coq. More recently, large language models (LLMs) have been used to generate complete proofs or sketches (jiang2022draft; cao2025reviving), achieving near-superhuman results on mathematical problems (trinh2024solving). Notably, AlphaProof (hubert2025olympiad) combines Monte Carlo tree search with reinforcement learning to solve IMO-level competition problems in Lean, representing the state of the art in search-based mathematical theorem proving.
Closer to our setting, prior efforts have attempted to bring such techniques to software verification, including Baldur (first2023baldur), PALM (lu2024proof), Guided Proof Search (prasad2024guided), and Goedel-Code-Prover (li2026goedel). However, these approaches remain limited to textbook-scale problems and do not address the challenges of real-world system-level verification projects such as seL4.
@@
Complementary to whole-proof generation, the proof-step prediction paradigm harnesses LLMs to propose the next tactic or lemma, guided by proof-state feedback. Systems such as GPT-f (polu2020generative), LISA (jiang2021lisa), HTPS (lample2022hypertree) and LeanDojo (yang2023leandojo) demonstrate the potential of this paradigm across Metamath, Isabelle and Lean, but again primarily on isolated mathematical problems. Our method fills this gap by integrating a fine-tuned LLM for proof search with symbolic techniques, namely proof-step revision and proof-state filtering, making large-scale verification both effective and efficient.
@@
To further enhance reliability, hybrid approaches are proposed to incorporate symbolic provers into LLM pipelines. A series of works, including Thor (jiang2022thor), ProofAug (liu2025proofaug), Strat2Rocq (fang2025proof), and LIPS (li2025proving), exemplify this by coupling LLM-generated reasoning with ATPs or hammers. However, hammer-only proofs tend to grow brittle on long, multi-step obligations and place additional trust in the external solver. In contrast, our framework integrates LLMs with an ITP-driven proof search and invokes ATP only as a final-stage closer for residual subgoals, keeping the bulk of the proof inside the kernel-checked workflow.
@@ 7 Limitations and Future Work
Rigid rule-driven step revision. The proof step revision modules heavily rely on heuristic tactic-premise recombination and lightweight premise retrieval. These mechanisms, though helpful and portable via statistical extraction, are imperfect and may introduce misleading candidates. Future improvements may include learning the relevance of premises directly from the proof state or integrating stronger semantic filters.
Leveraging frontier models. As frontier LLMs continue to advance in reasoning capabilities, an important direction is to explore whether agentic workflows, where a large model autonomously invokes verification tools, inspects proof states, and iterates, can complement or even replace parts of our pipeline. Our current use of fine-tuned small models is motivated by two practical advantages: lower cost and higher inference throughput, both critical for generating and scoring large candidate pools during tree search. A promising hybrid architecture would delegate high-level proof planning and subgoal decomposition to a frontier model, directly addressing the performance degradation on longer proofs discussed above, while relying on lightweight fine-tuned models for intensive step generation and ranking.
@@ 8 Conclusion
In this paper, we present a neuro-symbolic proof-generation framework that automates proof search for large-scale software verification projects in interactive theorem provers. Instead of attempting whole-proof synthesis, our approach performs a best-first tree search over fine-grained proof states, repeatedly querying a lightweight, fine-tuned LLM for the next proof step. This paradigm not only substantially improves data efficiency, but also enables a seamless integration of symbolic reasoning tools with LLM-based step generation and scoring, forming an effective and efficient pipeline for proof-step construction as well as proof-state filtering and ranking. The experiments on the seL4 verification corpus show that our approach significantly outperforms prior baselines. Notably, our method succeeds on a considerably higher proportion of multi-step, nontrivial theorems, addressing a long-standing limitation of LLM-based provers. Additional experiments on AFP projects and translated program-verification benchmarks further showcase strong cross-project generalizability.
Our work provides compelling evidence that proof-step–centric neuro-symbolic search can serve as a practical path toward scalable, automated software verification at the systems level, e.g., operating systems kernels, network architecture, and databases. While significant challenges remain, our results represent promising progress toward developing AI-enabled, rigorous engineering methods for real-world computer systems.
Availability. The tool, experimental data, and source code are available at the seL4-proof-search GitHub repository. Additionally, a prebuilt Docker image containing the seL4 verification environment and all required dependencies is available at Docker Hub.
Acknowledgments. We thank the anonymous reviewers and our shepherd for their valuable and constructive feedback. This work is supported by the Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (No. JYB2025XDXM118), the National Natural Science Foundation of China (Grant #62025202), and the Frontier Technologies R&D Program of Jiangsu (BF2024059). T. Chen is partially supported by overseas grants from the State Key Laboratory of Novel Software Technology, Nanjing University (KFKT2023A04, KFKT2025A05).